The Evanec Company's next expected dividend, D1, is $3.77; its growth rate is 6%; and its common stock now sells for $31.00. New stock (external equity) can be sold to net $24.80 per share.

  1. What is Evanec's cost of retained earnings, rs? Do not round intermediate calculations. Round your answer to two decimal places.

    rs =   %

  2. What is Evanec's percentage flotation cost, F? Round your answer to two decimal places.

    F =   %

  3. What is Evanec's cost of new common stock, re? Do not round intermediate calculations. Round your answer to two decimal places.

    re =   %
Answer #1

a) calculation of cost of retained earnings:

Cost of retained earnings= (D1/current price)+ growth

Cost of retained earnings= (3.77/31)+0.06= 0.1216+0.06= 0.1816

Cost of retained earnings= 18.16%

b) calculation of flotation cost percentage:

Floatation cost percentage= (31-24.80)/31*100= 20%

c) calculation of cost of new stock:

Cost of new stock= (3.77/24.80)+0.06= 0.1520+0.06= 0.2120

Cost of new stock is 21.20%